OpenStudy (anonymous):

A Chemist has a 20-mg sample of polonium-218. He needed approximately 81.5% of it for an experiment. given that the half-life of polonium-218 is approximately 3.1 min, how many seconds will it take for the sample to decay to the desired mass

OpenStudy (anonymous):

M=I(1.2)^t/n n = time of half-life

OpenStudy (kohai):

Hey, this is in the wrong section. You might receive quicker help if you post this into the science section.

OpenStudy (anonymous):

20x 0.815 = 16.3grams

OpenStudy (anonymous):

16.3= 20(1/2)^t/n

OpenStudy (anonymous):

16.3=20(1/2)^t/3.1 3.1 min =186 seconds